In Exercises 39-48, find a parametrization of the curve. 39. The vertical line passing through the point (3,2,0) 40. The line #48

#46 and #48
Answer #1

(46) = x² - y² and z= x² +ay-1 egnate both equations - x² - y2 = y²+ay-1 - 0 = y²tay- Plet y = t), > a = + -t) and ay = 1-t²There for parametersations [814)= < (7-7), t, (+2 -2) > S=1 center (3, 1,5) [x-3 Ixfixed and translated to have conter (3,11